Solve for the missing information designated by ? in the following table. (Use 365 days in a year. Round the in ratio to one decimal place before computing days to sell. Round days to sell to one decimal place.) ventory turnover Inventory Turnover Ratio Days to Sell Case Bl Purchases CGS El 150 $ 750 S 750 S 1,400 1,200 $ 150 5.0 7.0 73.0 250 C. 125 30.4
Answer #1

Answer

  • All working forms part of the answer
  • Answer

Case

BI

Purchases

CGS

EI

Inventory Turnover Ratio

Days to sell

a

$                                          150.00

$                 750.00

$                750.00

$              150.00

5.0

73.0

b

$                                          250.00

$             1,300.00

$            1,400.00

$              150.00

7.0

52.1

c

$                                             75.00

$             1,250.00

$            1,200.00

$              125.00

12.0

30.4

  • Working for above

Case

BI

Purchases

CGS

EI

Inventory Turnover Ratio

Days to sell

a

150

750

750

=150+750-750

=750/((150+150)/2)

=365/5

b

250

=1400+150-250

1400

=400-250

7

=365/7

c

=200-125

=1200+125-75

1200

125

=365/30.4

30.4

  • For Case ‘b’ first Average inventory is calculated as CGS / Inventory T Ratio = 1400 / 7 = 200

This means total of Bi and EI = 200 x 2 = 400

Then based on this EI and other are calculated as mentioned in above working table

  • For Case ‘c’, first Inventory T ratio is calculated as 365 days/30.4 = 12.0

The average inventory got calculated as CGS/IT Ratio = 1200/12 = 100

Total of BI and EI = 100 x 2 = $ 200